OpenStudy (anonymous):

Which equation demonstrates the Distributive Property? 3 • (8 + 2) = 3 • (2 + 8) 2 • (5 • 17) + 12 = (2 • 5) • 17 + 12 2 • (8 + 9) = 2 • 8 + 2 • 9 4 • (7 + 0) = 4 • 7

OpenStudy (anonymous):

a(b+c)=ab+ac

OpenStudy (anonymous):

2*(8+9)=2*8+2*9

OpenStudy (anonymous):

Si